OpenStudy (anonymous):

HELP!! multivariable calculus--> find the tangent plane to z=f(x,y) at the point P: f(x,y)=x^2*y-2x^3+y+3 with P=(-1,2,9)

OpenStudy (amistre64):

the partials will be needed for tangent planes

OpenStudy (amistre64):

can you solve for Fx and Fy?

OpenStudy (amistre64):

Fz is -1 if we allow: 0 = f(x,y) - z

OpenStudy (anonymous):

I feel lost on how to start this

OpenStudy (amistre64):

you need to know what partial derivatives are ....

OpenStudy (amistre64):

partials assume that anything other than the "wrt" variable is a constant. f(x,y)=x^2*y-2x^3+y+3 Fx, assume all other letters are constants Fx=2x*y -6x^2+0+0 --------------------------- Fy, assume all other letters are constants Fy=x^2 -0 +1+0

OpenStudy (amistre64):

this creates expressions for a gradient of f: (Fx,Fy,Fz), if memory serves, this also defines the normal of the tangent plane

OpenStudy (amistre64):

hence, given a point (a,b,c) Fx(x-a) + Fy(x-b) + Fz(z-c) = 0

OpenStudy (anonymous):

so how do you find what Fz is? and x,y,z

OpenStudy (amistre64):

xyz are just fillins for the plane equation; kinda like y = mx + b, asking what y and x are is futile.

OpenStudy (amistre64):

since z = f(x,y) 0 = f(x,y) - z we can say that g(x,y,z) = f(x,y) - z to rename things to find Gx, Gy, Gz with Gx = Fx Gy = Fy Gz = -1, which we can use to fill in the z spot on the normal vector

OpenStudy (anonymous):

is the answer z-9=-10(x+1)+2(y-2) ??

OpenStudy (amistre64):

maybe ... given a point: (-1,2,9) the plane is Fx(x+1) + Fy(y-2) + Fz(z-9) = 0 Fx = 2(-1)(2) -6(-1)^2 = -10 Fy = (-1)^2 +1 = 2 Fz = -1, just becuase -10(x+1) + 2(y-2) - (z-9) = 0 and since leading terms are somehow bad if they are negative .... 10(x+1) - 2(y-2) + (z-9) = 0 they also dont like it all bottled up like that and expect you to expand it all out. also, doble check my work to make sure i didnt do any simple mistakes :/